OpenStudy (anonymous):

Use Laplace transforms to solve the IVP: y''-2y'+5y=10 sin t y(0)=3, y'(0)=-4

OpenStudy (anonymous):

Thats what i have so far. \[(s^2-2s+5)Y-3s+10=\frac{ 10 }{ s^2+100 } \Rightarrow Y=\frac{ 3s^3-10s^2+200s-900 }{ (s^2+100)(s^2-2s+5) }\]

OpenStudy (anonymous):

\[\begin{align*} y''-2y'+5y&=10\sin t\\\\ \left(s^2Y-3s+4\right)-2\left(sY-3\right)+5Y&=\frac{10}{s^2+1}\\\\ \left(s^2-2s+5\right)Y&=\frac{10}{s^2+1}+3s-6\\\\ Y&=\frac{10+(3s-6)(s^2+1)}{(s^2+1)(s^2-2s+5)}\\\\ Y&=\frac{10+3s^3-6s^2+3s-6}{(s^2+1)(s^2-2s+5)}\\\\ Y&=\frac{3s^3-6s^2+3s+4}{(s^2+1)(s^2-2s+5)} \end{align*}\] Partial fractions: \[\begin{align*} \frac{3s^3-6s^2+3s+4}{(s^2+1)(s^2-2s+5)}&=\frac{As+B}{s^2+1}+\frac{Cs+D}{s^2-2s+5}\\\\ 3s^3-6s^2+3s+4&=(As+B)(s^2-2s+5)+(Cs+D)(s^2+1)\\\\ 3s^3-6s^2+3s+4&=(A+C)s^3+(-2A+B+D)s^2+(5A-2B+C)s\\&\quad+(5B+D) \end{align*}\] Find your constants: \[\begin{cases} \begin{align*}A+C&=3\\ -2A+B+D&=-6\\ 5A-2B+C&=3\\ 5B+D&=4\end{align*} \end{cases}\]

OpenStudy (anonymous):

@SithsAndGiggles Why do you have \[\frac{ 10 }{ s^2+1 } \] on the \[10 \sin t\]

OpenStudy (anonymous):

\[(s^2-2s+5)Y=\frac{ 10 }{ s^2+1 }+3s-10\] Thats what i got after i transformed your solution. You forgot about the -4

OpenStudy (anonymous):

Regarding your first reply: You're given \(10\sin t\), not \(\sin10t\). \[\mathcal{L}\{\sin at\}=\frac{a}{s^2+a^2}\]

OpenStudy (anonymous):

As for your second, you're right, sorry for the mix-up.

OpenStudy (anonymous):

That means you have \[Y=\frac{3s^3-10s^2+3s}{(s^2+1)(s^2-2s+5)}\] PFs:\[\begin{align*} \frac{3s^3-10s^2+3s}{(s^2+1)(s^2-2s+5)}&=\frac{As+B}{s^2+1}+\frac{Cs+D}{s^2-2s+5}\\\\ 3s^3-10s^2+3s&=(As+B)(s^2-2s+5)+(Cs+D)(s^2+1)\\\\ 3s^3-10s^2+3s&=(A+C)s^3+(-2A+B+D)s^2+(5A-2B+C)s\\&\quad+(5B+D) \end{align*}\] System: \[\begin{cases} \begin{align*}A+C&=3\\ -2A+B+D&=-10\\ 5A-2B+C&=3\\ 5B+D&=0\end{align*} \end{cases}\]